OpenStudy (anonymous):

find the 2nd Derivative y= 7x* sin(x^2) I tried to do it many times but i got my answer wrong but i dont know why ???

OpenStudy (anonymous):

Can you show your work so far? Should be pretty simple application of the product and chain rules.

OpenStudy (anonymous):

yeah i know i did the same and i got -2cosx^2-4x^2sinx

OpenStudy (anonymous):

How did you get that? You have f(x)=7x, g(x)=sin(x^2), product rule says you do f'(x)g(x)+g'(x)f(x).

OpenStudy (anonymous):

So we have f(x)=7x, g(x)=sin(x^2), f'(x)=7, g'(x)=2x cos(x^2), then just plug everything in.

OpenStudy (anonymous):

can you show me please so I can know what is my mistake ?

OpenStudy (anonymous):

What part haven't I already shown you? I've already broken it down into f(x) and g(x), and already calculated the derivatives of them, literally the only thing remaining to do is plug it back into the product rule formula.

OpenStudy (anonymous):

actually the 2nd derivative too not just the first one ?

OpenStudy (anonymous):

i found the first one now . 14x^2cosx^2+7sinx^2

OpenStudy (anonymous):

Okay, yes, that is correct for the first derivative. Now do the second derivative, again using the product and chain rules.

OpenStudy (anonymous):

second derivative is -28x^4sinx^2+14xcosx^2 is that right ?

OpenStudy (anonymous):

No, that is not correct. How did you get that, can you show your work?

OpenStudy (anonymous):

the correct answer is 42x cosx^2 -28x^3 sinx^2 im really confused with this problem i started doing it right but it turned at the end wrong ?

OpenStudy (anonymous):

Yes, I know that is the correct answer. You got the 14xcosx^2 part right, but you got the other part wrong. The product rule part. Can you show your work?

OpenStudy (anonymous):

-14x^2cosx^2*2x+7cos(x^2)*2x and i end with that answer

OpenStudy (anonymous):

Okay, no, that's not right at all. First step: split the product into f(x) and g(x). What are they?

OpenStudy (anonymous):

Remember the product rule? \[ (fg)'(x)=f'(x)g(x)+f(x)g'(x) \]

OpenStudy (anonymous):

i know what do you mean but i get anther problem and im working on it y= 4x sinx^2 f(x) 4x f'(x)= 4 g(x)= sinx^2 g'(x)= 2x cos(x^2) y'= 6x^2cosx^2+4sin x^2

OpenStudy (anonymous):

You mean 8, not 6, on that first term of y'

OpenStudy (anonymous):

yeah i missed it

OpenStudy (anonymous):

im doing now the second derivative

OpenStudy (anonymous):

i got -16x^3 sinx^2+8x cos x^2 can you check it ?

OpenStudy (anonymous):

You need to use the product rule to differentiate 6x^2cosx^2

OpenStudy (anonymous):

you mean 8x^2 cos x^2 i used product rule

OpenStudy (anonymous):

Yes, sorry, was copying what you had. You're missing a term. You got the fg' term but not the f'g term

OpenStudy (anonymous):

\[f'(x) = 7 \sin \left(x^2\right)+14 x^2 \cos \left(x^2\right)\\ f''(x)= 42 x \cos \left(x^2\right)-28 x^3 \sin \left(x^2\right) \]

OpenStudy (anonymous):

can you show me how did get the f"" ?

OpenStudy (anonymous):

I leave the details for you.

OpenStudy (anonymous):

but still your answer is wrong ? thats what the web homework said

OpenStudy (anonymous):

\[f'(x) = 7 \sin \left(x^2\right)+14 x^2 \cos \left(x^2\right)\\ f''(x) = 14 x \cos(x^2) + 28 x\cos(x^2) - 28 x^3 \sin(x^2)=\\42x\cos(x^2) - 28 x^3 \sin(x^2) \]

OpenStudy (anonymous):

Do you follow me?

OpenStudy (anonymous):

yeah but i dont know why it showed me wrong

OpenStudy (anonymous):

What did you get as an answer?

OpenStudy (anonymous):

24xcosx^2-16x^3sinx^2 this is what showed me

OpenStudy (anonymous):

3x cos(x^2) new one

OpenStudy (anonymous):

For a new question you should close the question and open a new one.

OpenStudy (anonymous):

For the second link, copy and paste in your browser

OpenStudy (anonymous):

ok thanks